OpenStudy (jenniferjuice):

Choose the equation that could be used to find two consecutive integers whose sum is 69.

OpenStudy (jenniferjuice):

n + (n + 2) = 69 n + (n + 1) = 69 n − (n + 1) = 69 n + 2n = 69

OpenStudy (mertsj):

What are consecutive integers? Can you give an example?

OpenStudy (jenniferjuice):

integers that are never ending i guess you could say

OpenStudy (mertsj):

That would be integers that are together when you are counting. Like 4 and 5 or 10 and 11

OpenStudy (jenniferjuice):

so um?

OpenStudy (mertsj):

So if one integer is x....what is the next integer?

OpenStudy (jenniferjuice):

69?

OpenStudy (jenniferjuice):

maybe ? >.< idk?!?!

OpenStudy (mertsj):

So 69 comes after x when you are counting?

OpenStudy (jenniferjuice):

noooo

OpenStudy (mertsj):

What does come after x when you are counting?

OpenStudy (jenniferjuice):

y?

OpenStudy (jenniferjuice):

1?

OpenStudy (mertsj):

Look at me

OpenStudy (mertsj):

The number that comes after 4 is 4+1 The number that comes after 10 is 10+1 The number that comes after 25 is 25+1

OpenStudy (jenniferjuice):